# Observational constraints on the Kerr and its several single-parameter modified spacetimes using quasi-periodic oscillation data

### v1

@@ V.1 Bayesian analysis
If $R>1$, it hints that $\bar{M}_{1}$ is superior to $\bar{M}_{2}$ in explaining the data, whereas if $R<1$, the opposite is true Bys-1 . The strength of evidence when comparing two models, $\bar{M}_{1}$ and $\bar{M}_{2}$, is evaluated using an empirical scale known as the “Jeffreys scale”. This scale, as described in Bys-3 ; Bys-4 ; Bys-5 , allows for model preference assessment based on the value of $|\ln R|$. Specifically, if $|\ln R|<1$, the evidence is deemed inconclusive. Weak evidence is indicated by values between $2.51\leq|\ln R|<$ 2.5, while moderate evidence is suggested by $52.5\leq|\ln R|<$ 5. Strong evidence corresponds to values in the range $105\leq|\ln R|<$ 10, and decisive evidence in favor of one model over the other is indicated by $|\ln R|\geq 10$. Below, we summarize the values of the Bayes factor $R$ and $|\ln R|$ for the nine types of modified geometries of Kerr relative to Kerr spacetime in Table 4.
@@ V.2 Analysis from Akaike Information Criterion
In this formula, $K$ represents the number of free parameters in the model, while $\mathcal{L}_{\max}$ denotes the maximum likelihood obtained using the optimal parameters $\Theta$. The term $-2\ln\mathcal{L}(\Theta\mid D)_{\max}$ in Equation is known as $\chi^{2}$, which assesses how well the model fits the data, while the term $2K$ in Equation reflects the complexity of the model AIC-5 ; AIC-6 ; AIC-7 ; AIC-8 ; AIC-9 .
@@ Appendix A Expressions for the frequencies of particle motion in Kerr spacetime and its nine modified versions
(I) Kerr spacetime:
@@
(II) Bardeen spacetime:
@@
(III) ABG spacetime:
@@
$+\left(2M^{2}\omega_{\phi}^{2}\left(r^{2}+Q_{1}^{2}\right)^{3/2}\left(r^{4}-7Q_{1}^{2}r^{2}-2Q_{1}^{4}\right)r^{4}-M\left(r^{2}+Q_{1}^{2}\right)\left(2\left(5r^{2}\omega_{\phi}^{2}+2\right)r^{8}+5\left(r^{2}\omega_{\phi}^{2}-1\right)Q_{1}^{2}r^{6}\right.\right.$
$\left.\left.-3\left(19r^{2}\omega_{\phi}^{2}+8\right)Q_{1}^{4}r^{4}-17\left(3r^{2}\omega_{\phi}^{2}+1\right)Q_{1}^{6}r^{2}-2\left(7r^{2}\omega_{\phi}^{2}+1\right)Q_{1}^{8}\right)+\sqrt{r^{2}+Q_{1}^{2}}\left(\omega_{\phi}^{2}r^{12}+\left(17r^{2}\omega_{\phi}^{2}+5\right)Q_{1}^{2}r^{8}\right.\right.$
@@
$+a+r^{2}\left(2M^{2}\left(r^{2}+Q_{1}^{2}\right)^{3/2}\left(r^{4}-7Q_{1}^{2}r^{2}-2Q_{1}^{4}\right)r^{2}+\sqrt{r^{2}+Q_{1}^{2}}\left(3\omega_{\phi}^{2}r^{12}+\left(23r^{2}\omega_{\phi}^{2}+3\right)Q_{1}^{2}r^{8}\right.\right.$
@@
$\left.-M\left(r^{2}+Q_{1}^{2}\right)\left(8\omega_{\phi}^{2}r^{10}+\left(26\omega_{\phi}^{2}Q_{1}^{2}+2\right)r^{8}+\left(30\omega_{\phi}^{2}Q_{1}^{4}-3Q_{1}^{2}\right)r^{6}+Q_{1}^{4}\left(14\omega_{\phi}^{2}Q_{1}^{2}-33\right)r^{4}\right.\right.$
@@
(IV) Hayward spacetime:
@@
(V) KN spacetime:
@@
$\omega_{\theta}^{2}=$ $r^{-6}\Big{[}-a^{4}\omega_{\phi}^{2}\left(Q_{2}-2Mr\right)+2a^{3}\omega_{\phi}\left(Q_{2}-2Mr\right)+a^{2}\left(4Mr^{3}\omega_{\phi}^{2}+2Mr-Q_{2}\left(2r^{2}\omega_{\phi}^{2}+1\right)+r^{4}\omega_{\phi}^{2}\right)$
$+2ar^{2}\omega_{\phi}\left(Q_{2}-2Mr\right)+r^{6}\omega_{\phi}^{2}\Big{]}.$ (A15)
(VI) KTN spacetime:
@@
(VII) BK spacetime:
@@
(VIII) Kerr-MOG spacetime:
@@
(IX) Kerr-Sen spacetime:
$\omega_{\phi}=r^{-6}\Big{[}-a^{4}\omega_{\phi}^{2}(Q_{3}-2Mr)+2a^{3}\omega_{\phi}(Q_{3}-2Mr)+a^{2}\left(-Q_{3}\left(2r^{2}\omega_{\phi}^{2}+1\right)+4Mr^{3}\omega_{\phi}^{2}+2Mr+r^{4}\omega_{\phi}^{2}\right)+2ar^{2}\omega_{\phi}(Q_{3}-2Mr)+r^{6}\omega_{\phi}^{2}\Big{]},$ (A25)
@@
(X) PFDM spacetime:

### v2

@@ V.1 Bayesian analysis
If $R>1$, it hints that $\bar{M}_{1}$ is superior to $\bar{M}_{2}$ in explaining the data, whereas if $R<1$, the opposite is true Bys-1 . The strength of evidence when comparing two models, $\bar{M}_{1}$ and $\bar{M}_{2}$, is evaluated using an empirical scale known as the “Jeffreys scale”. This scale, as described in Bys-3 ; Bys-4 ; Bys-5 , allows for model preference assessment based on the value of $|\ln R|$. Specifically, if $|\ln R|<1$, the evidence is deemed inconclusive. Weak evidence is indicated by values between $1\leq|\ln R|<2.5$, while moderate evidence is suggested by $2.5\leq|\ln R|<5$. Strong evidence corresponds to values in the range $5\leq|\ln R|<10$, and decisive evidence in favor of one model over the other is indicated by $|\ln R|\geq 10$. Below, we summarize the values of the Bayes factor $R$ and $|\ln R|$ for the nine types of modified geometries of Kerr relative to Kerr spacetime in Table 4.
@@ V.2 Analysis from Akaike Information Criterion
In this formula, $K$ represents the number of free parameters in the model, while $\mathcal{L}_{\max}$ denotes the maximum likelihood obtained using the optimal parameters $\Theta$. The term $2\ln\mathcal{L}(\Theta\mid D)_{\max}$ in Equation is known as $\chi^{2}$, which assesses how well the model fits the data, while the term $2K$ in Equation reflects the complexity of the model AIC-5 ; AIC-6 ; AIC-7 ; AIC-8 ; AIC-9 .
@@ Appendix A Expressions for the frequencies of particle motion in Kerr spacetime and its nine modified versions
(I) In the Kerr spacetime (2), the concrete formulas for three frequencies: $\omega_{\phi}$, $\omega_{r}$, and $\omega_{\theta}$ are, respectively
@@
In this paper $M$ and $a$ correspond to denote the mass and spin parameter of the BH, respectively.
(II) For a regular Bardeen BH (3), the expressions for $\omega_{\phi}$, $\omega_{r}$, and $\omega_{\theta}$ are as follows
@@
Here, the modification parameter $b$ represents the magnetic monopole charge.
(III) In the framework of general relativity coupled with nonlinear electrodynamics, one can get a singularity-free BH solution (4), called the ABG. With $Q_{1}$ denoting the charge, in this spacetime three frequencies can be derived as
@@
[…]
@@
[…]
@@
[…]
@@
(IV) For the BH solution described by the Hayward metric (5), a model parameter $l$ is introduced to quantify the deviation of the Hayward BH from the Kerr BH. In the Hayward spacetime, the expressions of $\omega_{\phi}$, $\omega_{r}$, and $\omega_{\theta}$ are expressed as
@@
(V) For the KN BH (6), a solution to the Einstein-Maxwell equations, with the charge parameter $Q_{2}$ the expressions for $\omega_{\phi}$, $\omega_{r}$, and $\omega_{\theta}$ can be provided by
@@
$\omega_{\theta}^{2}=$ $r^{-6}\Big{[}-a^{4}\omega_{\phi}^{2}\left(Q_{2}^{2}-2Mr\right)+2a^{3}\omega_{\phi}\left(Q_{2}^{2}-2Mr\right)+a^{2}\left(4Mr^{3}\omega_{\phi}^{2}+2Mr-Q_{2}^{2}\left(2r^{2}\omega_{\phi}^{2}+1\right)+r^{4}\omega_{\phi}^{2}\right)$
$+2ar^{2}\omega_{\phi}\left(Q_{2}^{2}-2Mr\right)+r^{6}\omega_{\phi}^{2}\Big{]}.$ (A15)
(VI) In the KTN solution (7), we have the forms for $\omega_{\phi}$, $\omega_{r}$, and $\omega_{\theta}$
@@
The gravitational magnetic monopole parameter $n$ denotes a unique dual mass effect.
(VII) In the BK spacetime (8), we have the forms for $\omega_{\phi}$, $\omega_{r}$, and $\omega_{\theta}$,
@@
Here the ”tidal charge” parameter $q$ encodes the effect of the extra dimensions on the gravitational field.
(VIII) In the Kerr-MOG spacetime (9), the modification parameter is defined as $\alpha=\frac{G_{K}}{G}-1$ with $G_{K}$ denoting the enhanced gravitational constant. And the frequencies $\omega_{\phi}$, $\omega_{r}$, and $\omega_{\theta}$ have the forms, respectively
@@
(IX) In the Kerr-Sen spacetime (10), the explicit expressions for $\omega_{\phi}$, $\omega_{r}$, and $\omega_{\theta}$ can be written as
$\omega_{\phi}=-\frac{M\left(\sqrt{2}\left(Mr+Q_{3}^{2}\right)^{2}\sqrt{\frac{M^{2}\left(2Mr+Q_{3}^{2}\right)}{\left(Mr+Q_{3}^{2}\right)^{2}}}-2aM^{3}\right)}{2a^{2}M^{4}-\left(Mr+Q_{3}^{2}\right)^{2}\left(2Mr+Q_{3}^{2}\right)},$ (A25)
@@
The modification parameter is written as $Q_{3}=Q^{2}/M$, where $Q$ denotes the electric charge.
(X) In the PFDM spacetime (11), the expressions for $\omega_{\phi}$, $\omega_{r}$, and $\omega_{\theta}$ are
@@
The modification parameter $k$ characterizes the intensity of dark matter.